OpenStudy (anonymous):

The formula for the surface area of a rectangular prism is a=2(lw+lh+wh) . Solve the formula for w

OpenStudy (anonymous):

lw+lh+wh=a/2 lw=a/2-lh-wh w=la/2-h-wh/l

OpenStudy (anonymous):

The answer is a-2lh ------ = w 2l+2h

OpenStudy (anonymous):

I just don't know how to get to that answer

OpenStudy (radar):